# There Was Never a Bottleneck in Concept Bottleneck Models

### v2

@@ 4 Experiments
In this section, we present a series of experiments designed to empirically demonstrate that CBMs fail to impose an effective bottleneck, even in simple settings. We also examine the consequences of this limitation. In contrast, we show that MCBMs successfully enforce a bottleneck, thereby mitigating the issues that arise in its absence. Further implementation details—including encoder architectures and training hyperparameters for each experiment—are provided in Appendix E and at https://github.com/antonioalmudevar/minimal_cbm.
@@ 4.1 Do CBMs and MCBMs Leak Information?
We next examine whether $\bm{n}_{y}$ and $\bm{n}_{\bar{y}}$ are present in the representations across different CBM variants and datasets. For this purpose, we define the following task–concept configurations: (i) MPI3D[29], where $\bm{y}$ is the object shape, $\bm{n}_{y}$ the horizontal axis, $\bm{n}_{\bar{y}}$ the vertical axis, and $\bm{c}$ the remaining generative factors; (ii) Shapes3D[30], where $\bm{y}$ is the shape, $\bm{n}_{y}$ the floor color and wall color, $\bm{n}_{\bar{y}}$ the orientation, and $\bm{c}$ the remaining factors; (iii) CIFAR-10[31], where $\bm{y}$ is the standard classification task, $\bm{c}$ consists of 64 of the 143 attributes extracted by Oikarinen et al. [32] using GPT-3 [33], and $\bm{n}_{y}$ the remaining attributes, since all nuisances are correlated with $\bm{y}$; and (iv) CUB[34], where $\bm{y}$ is the bird species, $\bm{c}$ includes concepts from twelve randomly selected attribute groups, and $\bm{n}_{y}$ the attributes from the remaining 20 groups. While MCBMs natively support multiclass concepts, the other baselines in our study are limited to binary concepts. To ensure fairness, factors with $k$ classes are therefore represented as $k$ binary concepts.
@@
MPI3D Shapes3D CIFAR-10 CUB
Vanilla $35.0\pm 1.9$ $45.5\pm 6.4$ $19.8\pm 0.7$ $3.8\pm 1.0$
CBM [9] $28.1\pm 0.5$ $18.1\pm 2.9$ $18.5\pm 0.7$ $3.8\pm 0.8$
CEM [19] $43.2\pm 5.2$ $15.8\pm 3.9$ $27.2\pm 0.8$ $3.9\pm 1.1$
ARCBM [18] $28.2\pm 1.7$ $18.4\pm 2.1$ $18.2\pm 0.6$ $3.9\pm 0.9$
SCBM [23] $24.3\pm 0.4$ $21.8\pm 1.4$ $18.3\pm 0.7$ $3.6\pm 0.8$
MCBM (low $\gamma$) $10.7\pm 0.1$ $2.5\pm 0.1$ $18.0\pm 0.5$ $3.4\pm 0.9$
MCBM (medium $\gamma$) $6.7\pm 0.2$ $0.2\pm 0.3$ $18.1\pm 0.5$ $2.8\pm 0.8$
MCBM (high $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$ $\bm{17.6\pm 0.5}$ $\bm{2.4\pm 1.0}$
Task-related information leakage
To measure the presence of a nuisance factor $n_{j}\in\bm{n}_{y}$ in $\bm{z}$, we estimate $I(N_{j};Z\mid C)$—the information about $n_{j}$ contained in $\bm{z}$ beyond what is explained by $\bm{c}$. Since this quantity is intractable, we approximate it as $\hat{I}(N_{j};Z|C)=\hat{H}(N_{j}|C)-\hat{H}(N_{j}|C,Z)$, where $\hat{H}(N_{j}\mid C)=-\sum_{k=1}^{N}\log h_{\psi}^{c}(\bm{c}^{(k)})$ and $\hat{H}(N_{j}\mid C,Z)=-\sum_{k=1}^{N}\log h_{\psi}^{cz}(\bm{c}^{(k)},\bm{z}^{(k)})$. Here, $h_{\psi}^{c}$ and $h_{\psi}^{cz}$ are MLP classifiers trained to predict $n_{j}$ from $\bm{c}$ and $(\bm{c},\bm{z})$, respectively. In Table 3, we report the average value of $\frac{\hat{I}(N_{j};Z\mid C)}{H(N_{j})}$ across all $n_{j}\in\bm{n}_{y}$, which we call Uncertainty Reduction Ratio (URR). From these results, we conclude that: (i) CBMs tend to reduce nuisance information compared to VMs, though not consistently; (ii) CEMs generally preserve the largest amount of nuisance information, often exceeding even VMs; (iii) ARCBMs and HCBMs show no systematic advantage over CBMs in terms of nuisance removal; and (iv) MCBMs provide the strongest reduction of nuisance information, particularly as $\gamma$ increases, which enforces a stricter bottleneck.
@@
Table 4: Average value of URR for task-unrelated nuisances.
@@
Neural networks typically discard input information, as their layers are non-invertible [35, 36]. Moreover, since $\bm{n}_{\bar{y}}$ is irrelevant for predicting $\bm{y}$, there is no incentive to retain it in the representation $\bm{z}$. One might therefore expect $\bm{z}$ to be free of $\bm{n}_{\bar{y}}$. However, prior work has shown that neural representations often preserve information not directly related to the task [24, 37]. To examine this, we analyze whether $\bm{n}_{\bar{y}}$ is present in $\bm{z}$ for MPI3D and Shapes3D—the only settings where $\bm{n}_{\bar{y}}$ is non-empty. Table 4 reports URR values for the nuisance variables in $\bm{n}_{\bar{y}}$ in. We find that: (i) as in Table 3, CEMs retain more nuisance information than even VMs; and (ii) MCBMs are the only models that consistently eliminate nuisances across all values of $\gamma$, as expected: with no incentive to preserve $\bm{n}_{\bar{y}}$ and an explicit penalty for doing so, such information is naturally discarded.
@@
CIFAR-10 CUB
CBM $84.8\pm 0.2$ $96.3\pm 0.1$
CEM $84.8\pm 0.2$ $96.3\pm 0.1$
ARCBM $84.3\pm 0.2$ $96.2\pm 0.1$
SCBM $84.3\pm 0.2$ $\bm{96.5\pm 0.1}$
MCBM (l $\gamma$) $\bm{84.9\pm 0.1}$ $96.3\pm 0.2$
MCBM (m $\gamma$) $\bm{84.9\pm 0.2}$ $96.1\pm 0.1$
MCBM (h $\gamma$) $84.8\pm 0.1$ $95.8\pm 0.3$
@@
MPI3D Shapes3D CIFAR-10 CUB
Vanilla $\pm 0.0 100.0±0.010$$\pm 0.0$ $2.1\pm 0.6 77$$7.4\pm 0.3$
CBM $99.9\pm 0.0$ $100.0\pm 0.0$ $3.8\pm 0.1 77$$7.6\pm 0.2$
CEM $\pm 0.0 100.0±0.010$$\pm 0.0$ $3.1\pm 0.3 77$$7.5\pm 0.3$
ARCBM $24.2\pm 0.6$ $29.7\pm 0.4$ $68.9\pm 0.3$ $75.5\pm 0.8$
SCBM $24.2\pm 0.5$ $29.7\pm 0.3$ $62.3\pm 0.0$ $73.5\pm 0.4$
MCBM (l $\gamma$) $92.7\pm 1.1$ $100.0\pm 0.0$ $72.4\pm 0.5$ $78.3\pm 0.6$
MCBM (m $\gamma$) $46.0\pm 0.6$ $32.2\pm 1.5$ $0.8\pm 0.2 77$$7.4\pm 0.4$
MCBM (h $\gamma$) $24.9\pm 0.3$ $30.1\pm 0.2$ $70.5\pm 0.6$ $73.5\pm 1.5$
As previously discussed, when $\bm{n}_{y}$ is non-empty, restricting $\bm{z}$ to contain only information about $\bm{c}$ should reduce performance on $\bm{y}$: if $\bm{c}$ alone is insufficient to solve $\bm{y}$, then $\bm{z}$ cannot achieve perfect accuracy. In Table 6, we observe the following: (i) CBMs and CEMs reach task accuracy comparable to (or even higher than) VMs, indicating that they also rely on $\bm{n}_{y}$ to predict $\bm{y}$; (ii) ARCBMs and SCBMs achieve lower task accuracy, as they impose a bottleneck after the representations (see Appendix D); and (iii) MCBMs show decreasing task accuracy as $\gamma$ increases, reflecting the stricter bottleneck applied to the representations. Importantly, this reduction in task accuracy should not be viewed negatively: since all models achieve similar concept accuracy (see Table 5), it indicates that predictions rely less on nuisance information.
@@ 4.2 Do MCBMs Yield More Interpretable Representations?
Figure 4: CKA (left) and Disentanglement (right).
One of the key properties often attributed to CBMs is that their internal representations align with human-interpretable concepts. While this generally holds, we show that MCBMs yield representations that are even more interpretable. To support this claim, we employ two metrics: (i) Centered Kernel Alignment(CKA) [38, 39, 40], which measures the similarity between learned representations and concept labels (encoded as one-hot vectors); and (ii) Disentanglement[41], which assesses whether each dimension of the representation $z_{j}$ captures at most one concept $c_{j}$. These metrics capture interpretability from complementary perspectives: (i) alignment indicates whether the information is organized in a concept-aware fashion; and (ii) disentanglement reflects the extent to which individual concepts are independently encoded in the representations. As shown in Figure 4: (i) CBMs, CEMs, ARCBMs, and HCBMs achieve stronger alignment with concepts than Vanilla Models, yet they do not consistently improve disentanglement; and (ii) MCBMs consistently improve both alignment and disentanglement, particularly as $\gamma$ increases, reflecting the removal of additional nuisances. These results suggest that explicitly eliminating nuisances leads to more interpretable representations.
@@ 4.3 Do MCBMs Enable More Reliable Interventions?
Another key property often attributed to CBMs is their capacity to support interventions. Although standard CBM intervention methods suffer from theoretical limitations (see Section 5), they can still be applied in practice. Accordingly, for CBMs we follow the procedure in Section 2.3, for ARCBMs and SCBMs we use the procedure described in Appendix D, and for MCBMs we apply the method in Section 2.4. To evaluate intervention reliability across models, we adopt the standardized protocol of Koh et al. [9], which tracks how prediction error changes as the number of intervened concepts increases. Following Shin et al. [42], we intervene on the concepts predicted with the lowest confidence. Results for CIFAR-10 and CUB are shown in Figure 5 (note that interventions have minimal effect on MPI3D and Shapes3D, since—as shown in Table 6—the concepts are not strong predictors of the task). From these results, four main observations emerge: (i) CBMs may even increase error when multiple concepts are intervened—an effect of nuisance information leaking into the representation due to the absence of a proper bottleneck; (ii) ARCBMs and SCBMs mitigate this by applying an Information Bottleneck after the representations, which improves intervention effectiveness while leaving interpretability unchanged (Figure 4); (iii) in MCBMs, intervention gains remain largely invariant to $\gamma$, as indicated by the nearly parallel trends across its different values; and (iv) at low and medium $\gamma$, MCBMs deliver the strongest intervention performance, especially when fewer than 100% of the concepts are intervened, clearly outperforming ARCBMs and SCBMs.

### v3

@@ 4 Experiments
In this section, we present a series of experiments designed to empirically demonstrate that CBMs fail to impose an effective bottleneck, even in simple settings. We also examine the consequences of this limitation. In contrast, we show that MCBMs successfully enforce a bottleneck, thereby mitigating the issues that arise in its absence. Further implementation details—including encoder architectures and training hyperparameters for each experiment—are provided in Appendix F.
@@ 4.1 Do CBMs and MCBMs Leak Information?
We next examine whether $\bm{n}_{y}$ and $\bm{n}_{\bar{y}}$ are present in the representations across different CBM variants and datasets. For this purpose, we define the following task–concept configurations: (i) MPI3D(Gondal et al., 2019), where $\bm{y}$ is the object shape, $\bm{n}_{y}$ the horizontal axis, $\bm{n}_{\bar{y}}$ the vertical axis, and $\bm{c}$ the remaining generative factors; (ii) Shapes3D(Kim and Mnih, 2018), where $\bm{y}$ is the shape, $\bm{n}_{y}$ the floor color and wall color, $\bm{n}_{\bar{y}}$ the orientation, and $\bm{c}$ the remaining factors; (iii) CIFAR-10(Krizhevsky et al., 2009), where $\bm{y}$ is the standard classification task, $\bm{c}$ consists of 64 of the 143 attributes extracted by Oikarinen et al. (2023) using GPT-3 (Brown et al., 2020), and $\bm{n}_{y}$ the remaining attributes, since all nuisances are correlated with $\bm{y}$; (iv) CUB(Wah et al., 2011), where $\bm{y}$ is the bird species, $\bm{c}$ includes concepts from twelve randomly selected attribute groups, and $\bm{n}_{y}$ the attributes from the remaining 15 groups; and (v) AwA2(Xian et al., 2017), where $\bm{y}$ is the animal class, $\bm{c}$ consists of 20 of the 85 human-annotated attributes, and $\bm{n}_{y}$ the remaining attributes. While MCBMs natively support multiclass concepts, the other baselines in our study are limited to binary concepts. To ensure fairness, factors with $k$ classes are therefore represented as $k$ binary concepts.
Task-related information leakage
To measure the presence of a nuisance factor $n_{j}\in\bm{n}_{y}$ in $\bm{z}$, we estimate $I(N_{j};Z\mid C)$—the information about $n_{j}$ contained in $\bm{z}$ beyond what is explained by $\bm{c}$. Since this quantity is intractable, we approximate it as $\hat{I}(N_{j};Z|C)=\hat{H}(N_{j}|C)-\hat{H}(N_{j}|C,Z)$, where $\hat{H}(N_{j}\mid C)=-\sum_{k=1}^{N}\log h_{\psi}^{c}(\bm{c}^{(k)})$ and $\hat{H}(N_{j}\mid C,Z)=-\sum_{k=1}^{N}\log h_{\psi}^{cz}(\bm{c}^{(k)},\bm{z}^{(k)})$. Here, $h_{\psi}^{c}$ and $h_{\psi}^{cz}$ are MLP classifiers trained to predict $n_{j}$ from $\bm{c}$ and $(\bm{c},\bm{z})$, respectively. In Table 3, we report the average value of $\frac{\hat{I}(N_{j};Z\mid C)}{H(N_{j})}$ across all $n_{j}\in\bm{n}_{y}$, which we call Uncertainty Reduction Ratio (URR). From these results, we conclude that: (i) CBMs tend to reduce nuisance information compared to VMs, though not consistently; (ii) CEMs and ECBMs generally preserve the largest amount of nuisance information, often exceeding even VMs; (iii) ARCBMs and HCBMs show no systematic advantage over CBMs in terms of nuisance removal; and (iv) MCBMs provide the strongest reduction of nuisance information, particularly as $\gamma$ increases, which enforces a stricter bottleneck.
@@
MPI3D Shapes3D CIFAR-10 CUB AwA2
Vanilla $35.0\pm 1.9$ $45.5\pm 6.4$ $19.8\pm 0.7$ $3.8\pm 1.0$ $1.5\pm 0.3$
CBM $28.1\pm 0.5$ $18.1\pm 2.9$ $18.5\pm 0.7$ $3.8\pm 0.8$ $1.4\pm 0.3$
CEM $43.2\pm 5.2$ $15.8\pm 3.9$ $27.2\pm 0.8$ $3.9\pm 1.1$ $1.1\pm 0.5$
ECBM $25.2\pm 3.0$ $47.1\pm 3.7$ $18.1\pm 0.5$ $4.5\pm 1.0$ $1.1\pm 0.3$
ARCBM $28.2\pm 1.7$ $18.4\pm 2.1$ $18.2\pm 0.6$ $3.9\pm 0.9$ $1.6\pm 0.3$
SCBM $24.3\pm 0.4$ $21.8\pm 1.4$ $18.3\pm 0.7$ $3.6\pm 0.8$ $1.2\pm 0.2$
MCBM (low $\gamma$) $10.7\pm 0.1$ $2.5\pm 0.1$ $18.0\pm 0.5$ $3.4\pm 0.9$ $1.0\pm 0.4$
MCBM (medium $\gamma$) $6.7\pm 0.2$ $0.2\pm 0.3$ $18.1\pm 0.5$ $2.8\pm 0.8$ $0.9\pm 0.4$
MCBM (high $\gamma$) $\bm{0.0\pm 0.0}$ $\bm{0.0\pm 0.0}$ $\bm{17.6\pm 0.5}$ $\bm{2.4\pm 1.0}$ $0.7\pm 0.4$
@@
Table 4: Average value of URR for task- nuisances.
@@
Neural networks typically discard input information, as their layers are non-invertible (Tishby and Zaslavsky, 2015; Tschannen et al., 2019). Moreover, since $\bm{n}_{\bar{y}}$ is irrelevant for predicting $\bm{y}$, there is no incentive to retain it in the representation $\bm{z}$. One might therefore expect $\bm{z}$ to be free of $\bm{n}_{\bar{y}}$. However, prior work has shown that neural representations often preserve information not directly related to the task (Achille and Soatto, 2018a; Arjovsky et al., 2019). To examine this, we analyze whether $\bm{n}_{\bar{y}}$ is present in $\bm{z}$ for MPI3D and Shapes3D—the only settings where $\bm{n}_{\bar{y}}$ is non-empty. Table 4 reports URR values for the nuisance variables in $\bm{n}_{\bar{y}}$ in. We find that: (i) as in Table 3, CEMs and ECBMs can retain more nuisance information than even VMs; and (ii) MCBMs are the only models that consistently eliminate nuisances across all values of $\gamma$, as expected: with no incentive to preserve $\bm{n}_{\bar{y}}$ and an explicit penalty for doing so, such information is naturally discarded.
@@
CIFAR-10 CUB AwA2
CBM $84.8\pm 0.2$ $96.3\pm 0.1$ $98.1\pm 0.1$
CEM $84.8\pm 0.2$ $96.3\pm 0.1$ $98.0\pm 0.1$
ECBM $84.6\pm 0.2$ $96.1\pm 0.4$ $98.0\pm 0.1$
ARCBM $84.3\pm 0.2$ $96.2\pm 0.1$ $97.9\pm 0.1$
SCBM $84.3\pm 0.2$ $\bm{96.5\pm 0.1}$ $\bm{98.4\pm 0.1}$
MCBM (l $\gamma$) $\bm{84.9\pm 0.1}$ $96.3\pm 0.2$ $98.0\pm 0.2$
MCBM (m $\gamma$) $\bm{84.9\pm 0.2}$ $96.1\pm 0.1$ $97.9\pm 0.1$
MCBM (h $\gamma$) $84.8\pm 0.1$ $95.8\pm 0.3$ $97.6\pm 0.2$
@@
As previously discussed, when $\bm{n}_{y}$ is non-empty, restricting $\bm{z}$ to contain only information about $\bm{c}$ should reduce performance on $\bm{y}$: if $\bm{c}$ alone is insufficient to solve $\bm{y}$, then $\bm{z}$ cannot achieve perfect accuracy. In Table 6, we observe the following: (i) CBMs, CEMs and ECBMs reach task accuracy comparable to (or even higher than) VMs, indicating that they also rely on $\bm{n}_{y}$ to predict $\bm{y}$; (ii) ARCBMs and SCBMs achieve lower task accuracy, as they impose a bottleneck after the representations (see Appendix D); and (iii) MCBMs show decreasing task accuracy as $\gamma$ increases, reflecting the stricter bottleneck applied to the representations. Importantly, this reduction in task accuracy should not be viewed negatively: since all models achieve similar concept accuracy (see Table 5), it indicates that predictions rely less on nuisance information.
@@
MPI3D Shapes3D CIFAR-10 CUB AwA2
Vanilla $\pm 0.0 100.0±0.010$$\pm 0.0$ $2.1\pm 0.6 77$$7.4\pm 0.3$ $91.6\pm 0.4$
CBM $99.9\pm 0.0$ $100.0\pm 0.0$ $3.8\pm 0.1 77$$7.6\pm 0.2$ $91.1\pm 0.4$
CEM $\pm 0.0 100.0±0.010$$\pm 0.0$ $3.1\pm 0.3 77$$7.5\pm 0.3$ $90.7\pm 0.5$
ECBM $99.9\pm 0.0$ $100.0\pm 0.0$ $4.4\pm 0.4 77$$7.1\pm 0.5$ $92.1\pm 0.3$
ARCBM $24.2\pm 0.6$ $29.7\pm 0.4$ $68.9\pm 0.3$ $75.5\pm 0.8$ $89.9\pm 0.3$
SCBM $24.2\pm 0.5$ $29.7\pm 0.3$ $62.3\pm 0.0$ $73.5\pm 0.4$ $91.8\pm 0.1$
MCBM (l $\gamma$) $92.7\pm 1.1$ $100.0\pm 0.0$ $72.4\pm 0.5$ $78.3\pm 0.6$ $90.6\pm 0.4$
MCBM (m $\gamma$) $46.0\pm 0.6$ $32.2\pm 1.5$ $0.8\pm 0.2 77$$7.4\pm 0.4$ $90.1\pm 0.2$
MCBM (h $\gamma$) $24.9\pm 0.3$ $30.1\pm 0.2$ $70.5\pm 0.6$ $73.5\pm 1.5$ $88.7\pm 0.3$
@@ 4.2 Do MCBMs Yield More Interpretable Representations?
One of the key properties often attributed to CBMs is that their internal representations align with human-interpretable concepts (Debole et al., 2025). While this generally holds, we show that MCBMs yield representations that are even more interpretable. To support this claim, we employ two metrics: (i) Centered Kernel Alignment(CKA) (Cristianini et al., 2001; Cortes et al., 2012; Kornblith et al., 2019), which measures the similarity between learned representations and concept labels (encoded as one-hot vectors); (ii) Disentanglement(Eastwood and Williams, 2018), which assesses whether each dimension of the representation $z_{j}$ captures at most one concept $c_{j}$; and (iii) Oracle Information Score (OIS) (Zarlenga et al., 2023), which dextends beyond standard disentanglement by accounting for concept dependencies, requiring that correlations between $z_{j}$ and $z_{k}$ do not exceed those between $c_{j}$ and $c_{k}$. These metrics capture interpretability from complementary perspectives: (i) CKAindicates whether the information is organized in a concept-aware fashion; and (ii) Disentanglement and OIS reflect the extent to which individual concepts are independently encoded in the representations. As shown in Figure 4: As shown in Figure 4: (i) CBMs, CEMs, AR-CBMs, and HCBMs generally achieve higher CKA with concepts than Vanilla Models, but they do not consistently improve Disentanglement or OIS; (ii) ECBMs do not, in general, produce more interpretable representations than VMs; and (iii) MCBMs consistently improve CKA, Disentanglement, and OIS, with larger gains as $\gamma$ increases, reflecting the removal of additional nuisances.
These results indicate that explicitly removing nuisances leads to more interpretable representations. To analyze this systematically, Table 7 reports the rank correlations between the metrics in Figure 4 and the URR values in Table 3 across datasets. We observe that CKA and OIS are strong and consistent predictors of nuisance information, whereas disentanglement is noticeably weaker. This is particularly helpful for tuning $\gamma$: since $n_{y}$ is unobserved in real scenarios, URR cannot be computed as in Table 3. As a workaround, these metrics—computed solely from $\bm{c}$ and $\bm{z}$—can serve as practical proxies for selecting the trade-off between task accuracy and leakage.
Figure 4: CKA (left, ↑), Disentanglement (middle, ↑) and OIS (right, ↓).
Table 7: Rank Correlation (p-value) between each metric and URR for $n_{y}$ across datasets
MPI3D Shapes3D CIFAR-10 CUB AwA2
CKA .95(<.001)-.95\,(<\!.001) .91(<.001)-.91\,(<\!.001) .71(.04)-.71\,(.04) .77(.02)-.77\,(.02) .75(.04)-.75\,(.04)
Disentanglement .81(.01)-.81\,(.01) .57(.13)-.57\,(.13) .43(.29)-.43\,(.29) .69(.05)-.69\,(.05) .99(<.001)-.99\,(<\!.001)
OIS .90(.002).90\,(.002) .81(.01).81\,(.01) .86(.006).86\,(.006) .89(.002).89\,(.002) .66(.07).66\,(.07)
@@ 4.3 Do MCBMs Enable More Reliable Interventions?
Another key property often attributed to CBMs is their capacity to support interventions. Although standard CBM intervention methods suffer from theoretical limitations (see Section 5), they can still be applied in practice. Accordingly, for CBMs we follow the procedure in Section 2.3, for ARCBMs and SCBMs we use the procedure described in Appendix D, and for MCBMs we apply the method in Section 2.4. To evaluate intervention reliability across models, we adopt the standardized protocol of Koh et al. (2020), which tracks how prediction error changes as the number of intervened concepts increases. Following Shin et al. (2023), we examine the effect of interventions using two complementary policies on CIFAR-10, CUB, and AwA2: (i) intervening on the concepts with the lowest predicted confidence (results in Figure 5); and (ii) intervening on randomly selected concepts (results in Appendix E). For MPI3D and Shapes3D, interventions yield only minor changes in performance, which aligns with Table 6 showing that their concepts are relatively weak predictors of the target. Across datasets where concepts are informative, both intervention policies lead to consistent and converging conclusions: (i) CBMs may even increase error when multiple concepts are intervened—an effect of nuisance information leaking into the representation due to the absence of a proper bottleneck; (ii) ARCBMs and SCBMs mitigate this by applying an Information Bottleneck after the representations, which improves intervention effectiveness while leaving interpretability unchanged (Figure 4); (iii) in MCBMs, intervention gains remain mostly insensitive to $\gamma$ when only a small or moderate number of concepts are intervened—evidenced by the nearly parallel curves across $\gamma$ values—but increase with higher $\gamma$ when more concepts are intervened, as reflected in the steeper negative slopes; (iv) at low and medium $\gamma$, MCBMs deliver the strongest intervention performance, especially when fewer than 100% of the concepts are intervened, clearly outperforming ARCBMs and SCBMs; and (v) for AwA2—where all models exhibit low levels of nuisance retention (Table 3)—intervention performance is nearly identical across methods.